OpenStudy (anonymous):

k+4 and 2/3=9 and 3/4. What does k equal???

OpenStudy (anonymous):

Still don't get it... :P

OpenStudy (anonymous):

what don't you get about it?

OpenStudy (anonymous):

How can you find k?

OpenStudy (anonymous):

inverse operations

OpenStudy (anonymous):

I don't understand what your 2/3 =9 and 3/4 stand for either.

OpenStudy (anonymous):

Mind showing me one?

OpenStudy (anonymous):

show you one what???

OpenStudy (anonymous):

One of these problems. How you solve them.

OpenStudy (anonymous):



OpenStudy (anonymous):

Oh... why didn't you write it like that...

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Hey. I'll help you solve it.

OpenStudy (anonymous):

thank you!!!! help much needed. :P

OpenStudy (anonymous):

First, make the fractions into decimals so 9 2/3 = 9.666666666 and 9 3/4 is 9.75 k + 9.6666666 = 9.75 k + (9.66666666-9.6666666) = (9.75-9.66666666) k = 0.083333333

OpenStudy (anonymous):

Although you can write it like this:

OpenStudy (anonymous):



OpenStudy (anonymous):

it was 4 and 2/3 not 9 and 2/3

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

It's 5.083333 then

OpenStudy (anonymous):

thank you for the help you helped us a ton. :D :D :D

OpenStudy (anonymous):

No Problem.